Which firm has the highest EBITDA margin?

A

B

C

D

Sales

960

318

1550

725

COGS

730

233

1277

668

SG&A

72

35

90

49

Depreciation + Amort

165

28

60

27

EBIT

-7

22

123

-19

Net Income

-5

17

80

2
Answer #1

EBITDA=EBIT + Depreciation + Amortisation
EBITDA margin =EBITDA/Sales

For firm A:
EBITDA=-7 + 165=158
EBITDA margin =158/960=0.164583333


For firm B:
EBITDA=22 + 28=50
EBITDA margin =50/318=0.157232704


For firm C:
EBITDA=123 + 60=183
EBITDA margin =183/1550=0.118064516


For firm D:
EBITDA=-19 + 27=8
EBITDA margin =8/725=0.011034483

So, firm A has the highest EBITDA margin